OpenStudy (anonymous):

9x-2y=11 5x-2y=15 what is the solution to the above system of equations?

OpenStudy (anonymous):

You can convert them to y first, but because it is convinient, we'll change them to 2y: 9x-11=2y 5x-15=2y Then set the equations equal to each other(canceling out the 2y's): 9x-11=5x-15 Solve for x and then substitute it back into an equation and find y. Then, you'll have a coordinate (x,y) as your answer.

OpenStudy (anonymous):

Thank you so much

OpenStudy (anonymous):

This helped me so much thank you for showing it in such a simple and easy way @grazes

OpenStudy (anonymous):

No problem.

OpenStudy (anonymous):

I'm just surprised that you found this old question.